OpenStudy (anonymous):

Suppose that 4.5% of the people in the world have a particular genetic defect and that a screening test is 81% accurate for people who have it and 88% accurate for people who do not. If 1300 people are screened for the defect, which is the best prediction for the number of people with the defect who are identified as not having it? a. 11 b. 16 c. 48 d. 59 Please explain too! I don't get it!!!

OpenStudy (anonymous):

I am soooo confused... Thanks for coming to help me!!! Greatly appreciated :)

OpenStudy (anonymous):

the answer to that equation is 51.48... what do i do next @ParthKohli ??

OpenStudy (anonymous):

Answer should be 11

OpenStudy (anonymous):

The answer is 59 then????

OpenStudy (anonymous):

4.5% of the people have the genetic defect. Now consider,, we are screening 1300 people. So, 1300*4.5/100 ---> 58.5 has the defect. Again, screening test is successful 81% when the subject has defect. which means, 19% of the subjects having defect when screened are not diagnosed so out of 58.5, we have 58.5*19/100 is 11.11 Rounding to nearest integer, we have 11. Answer should be 11

OpenStudy (anonymous):

cuz 0.88*1300=58.5 which is about 59????

OpenStudy (anonymous):

oh i see @sumathy is it cuz 58.5 is the number of ppl who HAVE it and the remaining part (19%) is the ppl who DON"T have it?? so the answer is 11 then?!?!

OpenStudy (anonymous):

58.5 is the number of people who have it. But when they are screened, for 81% of them are diagnosed with. But 19% are wrongly identified to not having the defect.

OpenStudy (anonymous):

So the answer is 11 then???

OpenStudy (anonymous):

So the answer is 11 then???

OpenStudy (anonymous):

1300 -- total people (its been given that 4.5% generally have defect) So out of 1300 we can predict that 1300*4.5/100 =58.5 have the defect and other 1300*95.5/100 have no defect. As asked in the question "which is the best prediction for the number of people with the defect who are identified as not having it?" also given that when 100 people with defect go for the screening test 81% were rightly diagnosed with the defect. 19% thought they have the defect, they were wrongly diagnosed not to have it. so as per the question we need to find out the best prediction for the number of people having the defect but diagnosed wrongly not to have it. 58.5*19/100 =11

OpenStudy (anonymous):

Sorry, I just tried to explain one more time, but got myself more organized with second explanation :)

OpenStudy (anonymous):

@iheartfood yes answer is 11

OpenStudy (anonymous):

kk cool! thanks :) great explanation! i totally understand it now :)

OpenStudy (anonymous):

kk cool! thanks :) great explanation! i totally understand it now :)